a rectangle has the length that is two thirds its width. If the perimeter of the rectangle is 40 meters, what is the length of the rectangle?
Answer by lmeeks54(111)

Let L = the length
Let W = the width
Let P = the perimeter
...
The formula for the perimeter of a rectangle is P = 2L + 2W
...
Given:
P = 40 meters
L = 2/3W
...
Let's substitute the values for P and L into our formula for the perimeter and solve for W. After that, we can solve for L.
...
P = 2L + 2W = 40 m
2(2/3W) + 2W = 40
4/3W + 2W = 40
10/3W = 40
...
Multiply both sides by 3/10:
10/3W(3/10) = 40(3/10)
W = 12 m
...
L = 2/3W
L = 2/3(12)
L = 8 m
....
The answer to the question is L = 8 meters; however, it is best to always check your work.
...
P = 40 = 2(8) + 2(12)
40 = 16 + 24
40 = 40 checks
...
cheers,
Lee